\begin{thebibliography}{99}

\bibitem{Aschbacher}
Michael Aschbacher, Finite group theory. Cambridge Studies in Advanced Mathematics, 10. Cambridge
University Press, Cambridge, 1986.
\bibitem{Bender} H. Bender, Transitive Gruppen gerader Ordnung, in denen jede Involution genau einen Punkt fest l\"a{\ss}t, J. Algebra 17 (1971), 527 - 554.
\bibitem{Gorenstein} Daniel Gorenstein,  Finite groups. Harper \& Row, Publishers, New York-London 1968.
\bibitem{GLS2}Daniel  Gorenstein, Richard   Lyons and  Ronald
Solomon, The classification of the finite simple groups. Number 2. Part I.
 Chapter G.
General group theory. Mathematical Surveys and Monographs, 40.2. American Mathematical Society,
Providence, RI,  1996.

\bibitem{GLS3}Daniel  Gorenstein, Richard   Lyons and  Ronald
Solomon, The classification of the finite simple groups. Number 3. Part I.
 Chapter A.
Almost simple $K$-groups. Mathematical Surveys and Monographs, 40.3. American Mathematical Society,
Providence, RI,  1998.

\bibitem{huppert} B.  Huppert, Endliche Gruppen. I. Die Grundlehren der Mathematischen Wissenschaften,
Band 134 Springer-Verlag, Berlin-New York  1967 xii+793 pp.

\bibitem{MSS}Ulrich Meierfrankenfeld, Ulrich, Bernd Stellmacher and Gernot Stroth, Finite groups of local characteristic $p$: an overview.  Groups, combinatorics \& geometry (Durham, 2001),  155--192, World Sci. Publishing, River Edge, NJ, 2003.
\bibitem{WBN}  Christopher Parker and Peter Rowley,  Local characteristic $p$ completions of weak $BN$-pairs.
 Proc. London Math. Soc. (3)  93  (2006),  no. 2, 325--394.

\bibitem{ParkerStroth} Chris Parker and Gernot Stroth,  Strongly $p$-embedded subgroups, http://arxiv.org/abs/0901.0805, 2009.
\bibitem{Suzuki} M. Suzuki, On a class of doubly transitive groups II, Ann. Math. 79, 1964, 514 - 589.
\end{thebibliography}

\documentclass[12pt]{amsart}
\usepackage{amsmath, amssymb}
\newfont {\cyr} {wncyr10}
\renewcommand{\labelenumi}{{(\roman{enumi})}}

\usepackage{amsfonts}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{setspace}
\usepackage{multicol}

\newtheorem{theorem}{Theorem}[section]
\newtheorem{lemma}[theorem]{Lemma}
\newtheorem{proposition}[theorem]{Proposition}

\newtheorem{hypothesis}[theorem]{Hypothesis}

\newcounter{claim}[theorem]

\newcounter{cclaim}[theorem]

\def \udot {{}^{\textstyle .}}

\newcommand{\F}{\mathrm{F}}\newcommand{\B}{\mathrm{B}}\newcommand{\M}{\mathcal{M}}

\newcommand{\Aut}{\mathrm{Aut}}

\newcommand{\Syl}{\mathrm{Syl}}\newcommand{\syl}{\mathrm{Syl}}

\newcommand{\GF}{\mathrm{GF}}

\newcommand{\SL}{\mathrm{SL}}

\newcommand{\PGL}{\mathrm{PGL}}
\newcommand{\PSL}{\mathrm{PSL}}\newcommand{\PSp}{\mathrm{PSp}}
\newcommand{\Sym}{\mathrm{Sym}}
\newcommand{\Alt}{\mathrm{Alt}}
\newcommand{\Dih}{\mathrm{Dih}}

\def \syl {\hbox {\rm Syl}}\def \Syl {\hbox {\rm Syl}}

\def \ov {\overline}

\def \Aut{ \mathrm {Aut}}

\def \Fi {\mbox {\rm Fi}}

\def \J{\mbox {\rm J}}

\def \B{\mbox {\rm B}}

\def \M{\mbox {\rm M}}

\def \McL{\mbox {\rm McL}}

\def \PSU {\mbox {\rm PSU}}

\begin{document}
\renewcommand{\labelenumi}{(\roman{enumi})}

\title  {On strongly $p$-embedded subgroups of Lie rank $2$}
 \author{Chris Parker}
  \author{Gernot Stroth}

\address{Chris Parker\\
School of Mathematics\\
University of Birmingham\\
Edgbaston\\
Birmingham B15 2TT\\
United Kingdom} \email{c.w.parker@bham.ac.uk}

\address{Gernot Stroth\\
Institut f\"ur Mathematik\\ Universit\"at Halle - Wittenberg\\
Theordor Lieser Str. 5\\ 06099 Halle\\ Germany}
\email{gernot.stroth@mathematik.uni-halle.de}

\email {}

\def \eps {\epsilon}
\def\diag {\mathrm {diag}}
\def\SO {\mathrm{SO}}\def\GO {\mathrm{GO}}
\date{\today}

\maketitle \pagestyle{myheadings}

\markright{{\sc }} \markleft{{\sc Chris Parker and Gernot Stroth}}

\section{Introduction}

For a prime $p$, a proper subgroup $H$ of the finite group $G$ is  \emph{strongly $p$-embedded in
$G$} if $p$ divides $|H|$  and $p$ does not divide $|H \cap H^g|$ for all $g \in G\setminus H$. A
characteristic  property  of strongly $p$-embedded subgroups is that  $N_G(X)\le H$ for any
non-trivial $p$-subgroup $X$ of $H$. Strongly $p$-embedded subgroups appear in the final stages of
one of the programmes to better understand the classification of the finite simple groups \cite
{MSS}. The almost simple groups  with strongly $2$-embedded subgroups were determined by Bender
\cite{Bender} and Suzuki \cite{Suzuki}.

Recall that a \emph{$\mathcal K$-group} is a group in which every composition factor is from the
list of ``known" simple groups. That is, every simple section is either a cyclic group of prime
order, an alternating group, a group of Lie type or one of the twenty six sporadic simple groups.
In the classification of groups of local characteristic $p$, certain normalizers of non-trivial
subgroups are assumed to be $\mathcal K$-groups.

The objective of this paper is to extend the results of \cite[Corollary 1.4]{ParkerStroth} to cover
some of the Lie type groups of Lie rank 2 defined in characteristic $p$. Further results related to
local characteristic $p$ identifications of rank $2$ Lie type groups can be found in the work of
Parker and Rowley \cite{WBN}. Our main result in this article is the following theorem.

\begin{theorem}\label{thma}
Suppose that $p$ is an odd prime, $G$ is a finite group and that $N_G(T)$ is a $\mathcal K$-group
for all non-trivial $2$-subgroups $T$ of $G$. If $H$ is a  strongly $p$-embedded subgroup of $G$,
then $F^*(H)$ is not a simple Lie type group of Lie rank $2$ defined in characteristic $p$ unless
perhaps $F^*(H) \cong \PSL_3(p)$.
\end{theorem}

Suppose that $G$ and $H$ are as in Theorem~\ref{thma}. Then the main theorems in
\cite{ParkerStroth} can be applied when the $p$-rank of $C_H(t)$, $m_p(C_H(t))$, is at least $2$
for every involution $t$ in $H$. For groups $H$ with $F^*(H)$ a simple Lie type group defined in
characteristic $p$ and Lie rank $2$, we use \cite[Proposition 9.1]{ParkerStroth} to see that if
$m_p(C_H(t))\le 1$ for some involution in $H$, then $F^*(H) \cong \PSp_4(p)$ or $\PSL_3(p)$. Thus,
because of \cite[Theorem 1.2]{ParkerStroth}, to prove Theorem~\ref{thma}, we  now have to prove the
following theorem.

\begin{theorem}\label{thmone} Suppose that $p$ is an odd prime, $G$ is a finite group, $N_G(T)$ is a $\mathcal K$-group
for all non-trivial $2$-subgroups $T$ of $G$ and $H \le G$. If $H$ is strongly $p$-embedded in $G$,
then $F^*(H) \not \cong \PSp_4(p)$.
\end{theorem}

The methods that we apply in this paper cannot be easily extended  to eliminate the possibility
that $F^*(H)\cong \PSL_3(p)$. This is because in $\PSp_4(p)$ there are involutions which have
centralizer of $p$ rank $2$ whereas in $\PSL_3(p)$ there are no such involutions. This means that
it is much more difficult to control the structure of the centralizer of an involution of $G$ in
the latter case. Indeed so far we have been unable to eliminate this configuration, though the
classification of the finite simple groups shows that it cannot occur.

In Section 2 we present the facts about $\PSp_4(p) \cong \Omega_5(p)$ and background results  from
\cite{ParkerStroth} about strongly $p$-embedded subgroups. We prove  Theorem~\ref{thmone} in
Section 3. Our notation is standard, as can be found in any of \cite{Aschbacher, Gorenstein,
huppert}. It is also consistent with that in \cite{ParkerStroth}.

\section{Preliminaries}

In this section we present facts about $\PSp_4(p) \cong \Omega_5(p)$ and present background lemmas
which we require for the proof of Theorem~\ref{thmone}.

\begin{lemma}\label{O5facts} Suppose that $(V,Q)$ is a $5$-dimensional non-degenerate orthogonal space over  $\GF(p)$ where $p$ is  a prime. Let  $X =  \SO(V,Q) \cong \SO_5(p)$ and $Y \cong \Omega(V,Q)
 \cong \Omega_5(p)$.
Assume $\beta =\{e_1,e_2,e_3,e_4,e_5\}$ is an orthonormal basis for $(V,Q)$ and set $\eps =
(-1)^{(p-1)/2}$.
\begin{enumerate}
\item $\Aut(Y) \cong X$.
\item $Y$ has exactly two conjugacy classes of involutions, one has representative
$b=\diag(-1,-1,1,1,1)$ with respect to $\beta$ and the other has representative $a= \diag
(1,-1,-1,-1,-1)$ with respect to $\beta$.
\item As quadratic spaces, $[V,b]$ has $\eps$-type  and $[V,a]$ has $+$-type.
\item $C_Y(b)$ has a normal subgroup $L$ isomorphic to $\Omega_3(p)$ such that $C_Y(b)/L \cong
\mathrm O_2^\eps(p) \cong \Dih (2(p-\eps))$.
\item $a$ is $2$-central, $O^{2}(C_X(a)) \cong \Omega_4^+(p)$, $C_X(a)/O^{2}(C_X(a))$ is a fours
group and $O_{p'}(C_X(t))$ is a $2$-group.
\item $C_Y(b)$ contains exactly four conjugacy classes of fours groups containing $b$. They have representatives $F_1 =
\langle b, \diag (1,-1,-1,1,1) \rangle$ which contains three $Y$-conjugates of $b$, $F_2 = \langle
b, a \rangle$ which contains  $b$ and two $Y$-conjugates of $a$, $F_3 = \langle b, \diag
(1,1,-1,-1,1) \rangle$ which contains two $Y$-conjugates of $b$ and a $Y$-conjugate of $a$
and, finally, \begin{eqnarray*} F_4 &=& \langle b, r_vr_w\mid v \in [V,a], w \in C_W(a), Q(v)=Q(w)\\
&&\;\;\;\;\;\;\;\;\;\;\;\;\;\mbox{ a non-square and } r_x \mbox{ a reflection for }x\in \{v,w\}
\rangle\end{eqnarray*} which contains three $Y$-conjugates of $b$.

\end{enumerate}
\end{lemma}

\begin{proof} Parts (i) and (ii) are well known and can be read from \cite[Theorem 2.5.12, Table
4.5.1]{GLS3}. We have that $[V,b]=\langle e_1,e_2\rangle$ and  $[V,a]=\langle
e_2,e_3,e_4,e_5\rangle$. The type of these spaces is $\eps = (-1)^{m(p-1)/2}$ where $m$ is the
dimension of the space. This gives (iii). In $Z= \mathrm O_5(p)$, $C_Z(b)$  stabilizes  $[V,b]$ and
is thus isomorphic to $\mathrm O_2^\eps(p) \times \mathrm O_3(p)$. Note that $X$ has index $2$ in
$Z$ and $Y$ has index $4$ in $Z$ with $Z/Y$ elementary abelian. We obtain $C_Y(b)$ by taking the
subgroup of $C_Z(b)$ which contains $\Omega_2^\eps(p) \times \Omega_3(p)$ together with the
elements of $C_Z(b)$ which have determinant and spinor norm both $-1$ when projected to both
$\mathrm O ^\eps_2(p)$ and to $\mathrm O_3(p)$. So (iv) holds. Part (v) follows in the same way as
(iv).

It is straight forward to calculate all the $C_Y(b)$-classes fours groups as indicated  in (vi) by
considering the action of the fours groups on $[V,b]$.
\end{proof}

We will need the following consequence of the Thompson Transfer Lemma \cite[15.16]{GLS2}.

\begin{lemma}\label{TTLG} Suppose that $G$ is a group, $L= O^2(L) \le G$, $T \in \Syl_2(L)$ and $S \in \Syl_2(G)$
with $S\ge T$. Assume that there exists a dihedral subgroup $D \le S$ such that $D\cap T=1 $, $TD =
S$ and $C \le D$ is cyclic of index $2$ in $D$. If the involutions of $D\setminus C$ are not
$G$-conjugate to any involution in $TC$ and the involution in $C$ is not $G$-conjugate to any
involution in $T$, then $G$ has a normal subgroup $K$ such that $\syl_2(L) \subseteq \Syl_2(K) $
and $G= KD$.
\end{lemma}

\begin{proof} Since $|S:TC|=2$, the Thompson Transfer Lemma implies that $G$ has a normal subgroup $G_0$ of
index $2$. Then $L= O^2(L) \le O^2(G) \le G_0$ and  $S_0 = S \cap G_0= TD \cap G_0 = T(D\cap G_0)$.
Let $D_0 = D\cap G_0$. Then either $D_0= C$, $D_0 $ is dihedral of order $|D|/2$ or $|D_0|=2$.
Suppose that $D_0= C$. Then, because $L= O^2(L) \le O^2(G)$ and because the involution of $C$ is
not $G$-conjugate to an element of $T$, the Thompson Transfer Lemma applies repeatedly to give a
normal subgroup $K$ of $G_0$ which contains $L$  and has  $\syl_2(L) \subseteq \Syl_2(K) $ and $G=
KD$. Thus the lemma holds in this case. If $D_0$ is dihedral, then we may apply induction to $G_0$
and obtain  the required result. Thus we may assume that $D_0$ has order $2$. But then again the
Thompson Transfer Lemma gives the result.
\end{proof}

We will call on various  results from \cite[Section 3]{ParkerStroth} and for the sake of clarity we
repeat their statements here.

\begin{lemma}\label{main1} Suppose that $G$ is a group, $p$ is a prime and $H$ is a strongly
$p$-embedded subgroup of $G$. Then the following statements hold.
\begin{enumerate}
\item If $K \le G$, $K \not \le H$ and $H \cap K$ has order divisible by $p$, then $H \cap
K$ is a strongly $p$-embedded subgroup of $K$;
\item $\syl_p(H \cap K) \subseteq \syl_p(K)$; and
\item if $m_p(H) \ge 2$, then $O_{p'}(G) = \bigcap H^G$.
\end{enumerate}
\end{lemma}

\begin{proof}See \cite[Lemma 3.2]{ParkerStroth}.
\end{proof}

\begin{proposition} \label{SE-p2} Suppose that $p$ is a prime, $X$ is a $\mathcal K$-group and
$K=F^*(X)$ is simple. Let  $P \in \syl_p(X)$ and $Q= P\cap K$. If $m_p(P) \ge 2$ and $X$ possesses
a strongly $p$-embedded subgroup, then $K$ has a strongly $p$-embedded subgroup  and  $p$ and $K$
are as follows.
\begin{enumerate}
\item $p$ is arbitrary, $a \ge 1$ and $K \cong \PSL_2(p^{a+1})$,
or $\PSU_3(p^a)$,  ${}^2\B_2(2^{2a+1})$ $(p=2)$ or ${}^2{\rm G}_2(3^{2a+1})$ $(p=3)$ and $X/K$ is a
$p'$-group.
\item  $p > 3$, $K \cong \Alt(2p)$ and
$|X/K| \le 2$.
 \item $(K,p)$ is one of the pairs:  $(\PSL_2(8),3)$, $(\PSL_3(4),3)$,
 $(\M_{11},3)$, $({}^2\B_2(32),5)$, $({}^2\F_4(2)^\prime,5)$, $(\McL,5)$, $(\Fi_{22},5)$,
 $(\J_4,11)$.
\end{enumerate}\end{proposition}

\begin{proof} See \cite [7.6.1]{GLS3}. \end{proof}

We let $\mathcal E$ be the set of pairs $(K,p)$ in  Proposition~\ref{SE-p2} (i), (ii) and (iii)
with $p$ odd.

\begin{lemma}\label{main2} Suppose that $G$ is a group, $p$ is a prime and $H$ is a strongly
$p$-embedded subgroup of $G$. Set $\ov G = G/O_{p'}(G)$ and assume  further that $\ov H \neq \ov
G$. Then
\begin{enumerate}
\item $\ov H$ is strongly $p$-embedded in $\ov G$;
\item $F^*(\ov G)$ is a non-abelian simple group; and
\item if $G$ is a $\mathcal K$-group and $m_p(G) \ge 2$, then $(F^*(\ov G),p) \in \mathcal E$.
\end{enumerate}
\end{lemma}

\begin{proof}See \cite[Lemma 3.3]{ParkerStroth}.
\end{proof}

The next Lemma says that groups with strongly $p$-embedded subgroups have good control of
$G$-fusion in certain circumstances.

\begin{lemma}\label{control1} Suppose that $p$ is an odd prime, $G$ is a  group and  $H$ is strongly $p$-embedded of  $G$. Assume that for all involutions $t \in H$,  $p$ divides
$|C_H(t)|$. Then for all involutions $t \in H$, $t^G \cap H = t^H$.
\end{lemma}

\begin{proof}See \cite[Lemma 3.5]{ParkerStroth}.\end{proof}

\section{The proof of Theorem~\ref{thmone}}

In this section we prove Theorem~\ref{thmone}. Assuming that it is false we, work under the
following hypothesis.

\begin{hypothesis} $H$ is strongly $p$-embedded in a finite group $G$ and
$H^* = F^*(H)= E(H) \cong \PSp_4(p)\cong \Omega_5(p)$ for an odd prime $p$. Furthermore, $N_G(T)$
is a $\mathcal K$-group for all non-trivial $2$-subgroups $T$ of $G$.
\end{hypothesis}

\begin{lemma}\label{E=} Either $H = H^*$ or $H \cong \SO_5(p)$.
\end{lemma}

\begin{proof} This follows from Lemma~\ref{O5facts} (i) as $F^*(H) \cong\Omega_5(p)$.
\end{proof}

Let $s$ and $t$ be representatives of the two conjugacy classes of involutions in $H^*$ and assume
that $t$ is $2$-central in $H$. Furthermore assume that $F_3 = \langle s,t \rangle$ is the centre
of a Sylow $2$-subgroup $T$  of $C_H(s)$. Notice that $F_3$ contains exactly one $H$-conjugate of
$t$ by Lemma~\ref{O5facts} (iv). By Lemmas~\ref{main1} (iii) and \ref{E=}, $O_{p'}(G) =1$. Thus, as
$H^*$ is simple and contains a Sylow $p$-subgroup of $G$, if $K$ is a normal subgroup of $G$, $K
\cap H \ge H^*$. Thus, because of Lemma~\ref{main1} (iii), we may suppose by induction that $G$ is
a simple group. Since we may also assume that $G$ is not a $\mathcal K$-group, $G$ does not possess
a strongly $2$-embedded subgroup. In particular, by \cite[17.13]{GLS2}, we have the following
lemma.

\begin{lemma}\label{not2embedded} There is an involution $r \in H$ such that $C_G(r) \not \le
H$.\qed
\end{lemma}

The next lemma is the key result of the article. It shows that we can be much more specific about
the involution class which has centralizer not contained in $H$.

\begin{lemma}\label{Ctnotin} $C_G(t) \not \le H$.
\end{lemma}

\begin{proof} Suppose that $C_G(t) \le H$ and let $S \in \Syl_2(C_H(t))$ with $S \ge T$. Then, as $t$ is
$2$-central in $H$, we have $S \in \Syl_2(H)$. Furthermore, $t$ is the unique involution in $Z(S)$
and so $N_G(S) \le H$. Thus $S \in \Syl_2(G)$. Since $H$ controls $G$-fusion of involutions in $H$
by Lemma~\ref{control1}, $S \in \Syl_2(G)$, $G= O^2(G)$ and there are involutions in
$\SO_5(p)\setminus \Omega_5(p)$, the Thompson Transfer Lemma implies that $H = H^*$. Hence
Lemma~\ref{not2embedded} implies $C_G(s) \not \le H$. Since $s$ and $t$ are not $G$-conjugate by
Lemma~\ref{control1}, $t$ is the unique $G$-conjugate of $t$ in $F_3=Z(T)$. It follows that $N_G(T)
\le C_G(t) \le H$. Hence  $T$ is a Sylow $2$-subgroup of $C_G(s)$. From Lemma~\ref{O5facts}(iv),
$C_H(s)$ has cyclic Sylow $p$-subgroups of order $p$. As $H$ is strongly $p$-embedded in $G$,
Lemma~\ref{main1} (ii) implies that $C_G(s)$ has Sylow $p$-subgroups of order $p$. Let $P \in
\syl_p(C_H(s))$ and set $L= \langle P^{C_H(s)}\rangle$ and $R=O_{p'}(C_G(s))$. Then
Lemma~\ref{O5facts} (iv) implies that $L \cong \Omega_3(p)\cong\PSL_2(p)$. Assume that $R \not \le
H$.  Let $K$ be a component of $C_G(s)$ such that $p$ divides $|K|$. Then $K \not \le R$ and so $K$
centralizes $R$. Hence $R \le C_G(P) \le H$ which is a contradiction. Therefore $E(C_G(s)) \le R$.
As $P$ is not normal in $L$, $O_p(C_G(s))=1$. Hence $F(C_G(s)) \le R$ and, so $F^*(C_G(s)) \le R$.
Set $U = O(R)$ and let $E$ be a fours group of $L$. Then, for $e \in E^\#$, we have $\langle
e,s\rangle$ is a fours subgroup of $C_H(s)$ which is conjugate to $F_3$ by Lemma~\ref{O5facts}
(vi). It follows that $es$ is $H$-conjugate to $t$. Hence $C_{U}(e) = C_{U}(es)\le C_G(es) \le H$,
by hypothesis. Since $U= \langle C_{U}(e)\mid e \in E^\#\rangle$, we conclude that $U  \le H$ and
that $[U,L]=1$. Note that $T \cap R \in \syl_2(R)$ and $T \cap R \le O_{2}(C_H(s))$. Assume that $p
\neq 3$. Then $O_2(C_H(s))$ commutes with $L$. Hence $O_2(C_G(s)) $ is also centralized by $L$.
Thus $L$ commutes with $UO_2(R)$ and consequently $L$ commutes with $F(C_G(s))$. It follows that
$E(C_G(s)) \neq 1$. Let $K$ be a component of $C_G(s)$. Then $T \cap K \in \syl_2(K)$ and is
centralized by $L$. It follows that $L$ normalizes $K$. Since $L$ is a non-abelian simple group,
the Schreier property of simple groups implies that $L$ induces inner automorphisms of $K$
(remember $C_G(s)$ is a $\mathcal K$-group). As $L$ has order divisible by  $p$ and $K$ does not,
$[K,L]=1$. But then $K \le C_G(P) \le H$ and we have a contradiction. Hence, if $p \neq 3$, then $R
\le H$. Suppose that $p=3$.  If $K$ is a component of $C_G(s)$, then, as $K \le R$, $K$ is a
$3'$-group. Hence, as $C_G(s)$ is a $\mathcal K$-group,
 $K/Z(K)$ is  a Suzuki group. This, however, contradicts $2^6\le |K|_2 \le
|C_G(s)|_2=|T|=2^5$. Hence $E(C_G(s)) = 1$. Since $U \le H$, $C_{H}(s)$ is a $\{2,3\}$-group and $R
\cap H$ is
 a $2$-group, we infer that $U=1$ and $F=F^*(C_G(s)) =O_2(C_G(s))\le T\le C_H(s)$. Now $O_2(C_H(s))$ is elementary
abelian of order $16$ and so we have $F= O_2(C_H(s))$. But  $F$ contains exactly $5$ conjugates of
$t$ and so, as $H$ controls $G$-fusion of involutions in $H$ and, from the subgroup structure  of
$\PSp_4(3)\cong \PSU_4(2)$, $N_H(F)/F \cong \Alt(5)$, we infer that $C_G(s) \le H$, which is a
contradiction. Thus $R \le H$.

As $C_G(s) > RC_H(s)$, it follows from Lemma~\ref{main2}(ii) that $C_G(s)/R$ is an almost simple
group and $C_H(s)/R$ is strongly $p$-embedded in $C_G(s)/R$.  As $H=H^*\cong \Omega_5(p)$, the
structure of $C_H(s)$ given in Lemma~\ref{O5facts} (iv) implies $C_H(s)/L \cong \mathrm
O_2^\eps(p)$ where $\eps = (-1)^{(p-1)}$ and the extension is split. In particular, $C_H(s)/L$ has
dihedral Sylow $2$-subgroups. Set $W = T\cap L$ and let $D \le T$ be a dihedral $2$-group with $D
\cap W=1$ and $T= DW$. Let $C \le D$ be cyclic of index $2$ in $D$. Furthermore, make these choices
so that $s \in C$. Then $T_0 = CW$ has index $2$ in $T$ and, by Lemmas~\ref{O5facts} and
\ref{control1} all the fours groups which contain $s$ and are contained in $T_0$ are
$C_G(s)$-conjugate to $F_3$ and the fours groups containing $s$ and not contained in $T_0$ are not
$C_G(s)$-conjugate to $F_3$. Thus, as the involution in $C$, namely $s$, is not $C_G(s)$-conjugate
to an element of $W$ and $L= O^2(L)$, Lemma~\ref{TTLG} implies that $C_G(s)$ contains a normal
subgroup $K$, with $\Syl_2(L) \subseteq \Syl_2(K)$ and $C_G(s) = DK$. In particular, $K$ has
dihedral Sylow $2$-subgroups.

Since $L$ is simple when $p \ge 5$ and when $p=3$, $L$ has abelian Sylow $2$-subgroups of order
$4$, it follows from \cite[Theorem 16.3 pg.462]{Gorenstein} that $K/O(K) \cong \PSL_2(r^a)$ for
some odd prime $r$ or $K/O(K)\cong \Alt(7)$. Suppose first that $K/O(K) \cong \Alt(7)$. Then as $L$
contains a Sylow $2$-subgroup of $K$, we infer that $L \cong \PSL_2(7)$ and that $p=7$. Since
$TLR/R \cong \SO_3(7) \cong \PGL_2(7)$ and $\PGL_2(7)$ has no subgroup of index $7$, we have a
contradiction. So suppose that $K/O(K) \cong \PSL_2(r^a)$. Then, as $\PSL_2(p) \cong L \le H\cap K$
and $H\cap K$ is strongly $p$-embedded in $K$, we deduce from Dickson's List of subgroups of
$\PSL_2(r^a)$ \cite[8.27, pg. 213]{huppert} that $r\neq p$ and that $p=5$ or $p=3$. In  both cases
$L$ has elementary abelian Sylow $2$-subgroups of order $4$ and, from the structure of $\PSp_4(3)$
and $\PSp_4(5)$,  $O_{p'}(C_{C_H(s)}(P))$ is also a four group, where we recall that  $P \in
\syl_p(L)$. It follows that $O(K)=1$ and that $P$ is self-centralizing in $K$. On the other hand,
we know that the centralizer of $P$ in $K$ has order $(r^a+1)/2$ or $(r^a-1)/2$ and $K$ has abelian
Sylow $2$-subgroups. Therefore we infer that either $p=5$ and $r^a =11$ or $p=3$ and $r^a=5$. If
$p=5$, then $DL/C_D(L) \cong \PGL_2(5)$. But $\PGL_2(11)$ contains no such subgroup and this is a
contradiction. Thus $p=3$. Hence $H \cong \PSp_4(3)$, $L \cong \PSL_2(3)\cong \Alt(4)$, $D \cong
\Dih(8)$ and $K \cong \PSL_2(5)\cong \Alt(5)$. We also have $DL/C_{DL}(L) \cong \Sym(4)$ and so
$DK/C_D(K)\cong \Sym(5)$ and $|C_{D}(K)| =2^2$ . Since $P$ centralizes $C_{D}(K) $ and
$C_{DK/C_D(K)}(PC_D(K)/C_D(K)) \cong \Sym(3) \times \Sym(2)$, we have $|C_{DK}(P)| =2^33$. This
shows that  $C_{DK}(P) \not \le H$, which is a contradiction. Therefore we have shown $C_G(t) \not
\le H$ as claimed.
\end{proof}

We may now conclude the proof of Theorem~\ref{thmone}.

\begin{proof}[Proof of Theorem~\ref{thmone}] Since $C_G(t) \not\le H$ by Lemma~\ref{Ctnotin} and $m_p(C_G(t)) =2$,
Lemmas~\ref{main1} (iii) and \ref{main2} imply that $O_{p'}(C_G(t))\le H$ and, as $C_G(t)$ is a
$\mathcal K$-group, $(F^*(C_G(t)/O_{p'}(C_G(t))),p)\in \mathcal E$.  In particular, we have
$O_{p'}(C_G(t)) \le O_{p'}(C_H(t)) = O_2(C_H(t))$ by Lemma~\ref{O5facts} (v).

Suppose that $p=3$. Then, as $C_H(t)$ acts irreducibly on $O_2(C_H(t))/\langle t \rangle$, either
$O_{p'}(C_G(t))=\langle t \rangle$ or  $O_{p'}(C_G(t))= O_2(C_H(t))$. If $O_{p'}(C_G(t))=\langle
t\rangle$, then  $|O_2(C_H(t)/\langle t\rangle)| = 16$. As $C_H(t)/O_{p'}(C_G(t))$ is strongly
$p$-embedded in $C_G(t)/O_{p'}(C_G(t))$, this contradicts the structure of the groups in
$\mathcal{E}$ given in Proposition~\ref{SE-p2}  (see also \cite[Proposition 2.7]{ParkerStroth}).
Thus $O_2(C_H(t)) = O_2(C_G(t))= F^\ast(C_G(t))$ is extraspecial of order $2^5$ and plus-type.
Therefore we have $C_G(t)/O_2(C_G(t))$ is isomorphic to a subgroup of $\mathrm O^+_4(2)$. This then
forces $C_G(t) = C_H(t)$, which is impossible. Hence $p\neq 3$.

So assume that $p \ge 5$. Then, by Lemma~\ref{O5facts} (iv), $O_{p'}(C_G(t)) =\langle t\rangle$ and
$C_H(t)/\langle t\rangle$ is not soluble. The  structure of $F^*(C_H(t))/\langle t \rangle$
together with Proposition~\ref{SE-p2} implies that $p=5$ and that $F^*(C_G(t))/O_{p'}(G_G(t)) \cong
\Alt(10)$. Since $F^*(C_H(t)) \ge \langle t \rangle$, it follows that $F^*(C_G(t)) \cong 2\udot
\Alt(10)$. Now from the structure of $2\udot\Alt(10)$, we deduce that $X=F^*(C_H(t)) \cong
\SL_2(5)*\SL_2(5)$ and $N_{F^*(C_G(t))}(X)/X$ is cyclic of order $4$. This is unfortunately
incompatible with the structure of $N_G(X)$ in $H$ where we have either $N_H(X)/X$ is cyclic of
order $2$ or is a fours group. This contradiction shows that $H$ cannot be strongly $p$-embedded in
$G$ and completes the proof of Theorem~\ref{thmone}.
\end{proof}

As mentioned in the introduction, Theorem~\ref{thma} follows by combining \cite[Proposition 9.1,
Theorem 1.2]{ParkerStroth} with Theorem~\ref{thmone}.

\end{document}